# ScaleBiO: Scalable Bilevel Optimization for LLM Data Reweighting

### v1

@@ Appendix A Additional Experiments
A.1 Detailed MT-Bench Scores for LLaMA
The detailed MT-Bench scores for LLaMA-3-8B in Section B.2 are provided in Table 4.
Model Coding Extraction Humanities Math Reasoning Roleplay STEM Writing Average
LLaMA-3-8B Baseline 4.65 6.20 7.50 3.40 5.25 7.70 7.00 7.20 6.11
Ours 5.40 8.55 9.68 3.45 5.30 8.20 8.73 7.63 7.12
Table 4: Comparison of our approach and the baseline on MT-Bench.
@@ B.1 Small Scale Experiments
Throughout our small-scale experiments, we use GPT-2 [Radford et al., 2019] with 124 million parameters as the backbone model. For bilevel optimization hyperparameters, we set the learning rate to $10^{-2}$ for sampling weights and $10^{-5}$ for models. We run our algorithm for 3 epochs with a batch size of 64 and alpha of 10 while adopting AdamW [Loshchilov and Hutter, 2017] for optimization.
@@ B.2 Large Scale Experiments
Datasets Size Kind License
AlpacaGPT4 [Peng et al., 2023] 52K Instruction Apache-2.0
ShareGPT4 [Chiang et al., 2023] 6K Conversation Apache-2.0
SlimOrca [Lian et al., 2023] 518K Instruction MIT
AlpacaChat [Bian et al., 2023] 20K Conversation Apache-2.0
OpenOrcaGPT4 [Mukherjee et al., 2023] 1M Instruction MIT
WildChat [Zhao et al., 2024] 1M Conversation AI2 ImpACT
LMSYS-Chat [Zheng et al., 2023b] 1M Conversation LMSYS-Chat-1M
GPTeacher ["Teknium", 2023] 89K Instruction MIT
Airoboros [Durbin, 2023] 59K Conversation CC-BY-4.0
@@
Table 5: In this table, we summarize the datasets we used. Above the middle line are the high-quality datasets we used. Below the middle line are multilingual Alpaca datasets with language code in their name.
@@ Appendix C Important Lemmas
The results of Lemma 2 can be found in [Kwon et al., 2023] and Lemma B.1 of [Chen et al., 2023]. From Lemma B.7 in [Chen et al., 2023], the following result holds for $\Gamma^{\alpha}(\lambda)$:
@@
The result in Lemma 4 follows from Lemma B.2 of Chen et al. [2023].
@@
    $w_{\ast}^{\alpha}(\lambda)$ is $\ell_{w_{\ast},0}$-Lipschitz continuous where $\ell_{w_{\ast},0}=3\kappa$.
@@
Claim (i) in Lemma 5 can be found in Lemma 2.2 of [Ghadimi and Wang, 2018] and Claim (ii) implies from Lemma 3.2 (setting $\lambda_{1}=\lambda_{2}$) of [Kwon et al., 2023].
@@
Under Assumption 1, if $\alpha>2\ell_{11}/\mu_{2}$, then $u_{\ast}(\lambda)$ is $\ell_{\nabla u_{\ast}}$-smooth where $\ell_{\nabla u_{\ast}}=\mathcal{O}\left(\frac{\kappa^{2}}{\mu_{2}}\left(\ell_{21}+1\right)\right)$ where the condition number $\kappa=\max\left\{\ell_{10},\ell_{11},\ell_{21},\ell_{22}\right\}/\mu_{2}$
Following Lemma A.3 of [Kwon et al., 2023] and recalling the Lipschitz continuous property of $u_{\ast}(\lambda)$ from Lemma 5, we have this claim is correct.
@@ Appendix D Proofs of Theorem 1
$\mathcal{L}_{D_{k}}^{\alpha}(\lambda,w,u):=L_{1}(\lambda,w;D_{\text{val}}^{k})+\alpha\left(L_{2}(\lambda,w;D_{\text{tr}}^{k})-L_{2}(\lambda,u;D_{\text{tr}}^{k})\right)$ (15)
where $D_{k}=\left\{D_{\text{tr}}^{k},D_{\text{val}}^{k}\right\}$ and $D_{\text{tr}}^{k},D_{\text{val}}^{k}$ are i.i.d. mini-batch samples from the training and validation datasets $D_{\text{tr}},D_{\text{val}}$, respectively. We use $\mathcal{F}_{k}$ to denote the random information before the iteration $(\lambda_{k},w_{k},u_{k})$, that is $\mathcal{F}_{k}:=\sigma\left(\left\{(\lambda_{k},\omega_{k},u_{k}),D_{k-1},\cdots,D_{1}\right\}\right)$. We use $\mathcal{C}_{k}=\sigma\left(\left\{j_{1},j_{2}\cdots,j_{t-1};r_{1},r_{2},\cdots,r_{t-1}\right\}\right)$ to denote the random information of variables $u,w$ for the randomized block coordinates before the iteration $k$.
@@
$+2(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}\left\|\nabla\Gamma^{\alpha}(\lambda_{k})\right\|^{2}+(1+1/\gamma_{1})\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}.$ (32)
@@
$\mathbb{E}\left[\left\|\omega_{\ast}^{\alpha}(\lambda_{k})-\omega_{k+1}\right\|^{2}\mid\mathcal{F}_{k},\mathcal{C}_{k}\right]\leq\left(1-\frac{\alpha\mu_{2}\eta_{w}}{2J}\right)r_{k}+\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}$ (33)
@@
$R_{k}=\Gamma^{\alpha}(\lambda_{k})-\Gamma_{\min}^{\alpha}+\xi_{1,k}\delta_{k}+\xi_{2}^{k}r_{k}$ (35)
where $\xi_{1,k},\xi_{2,k}>0$ are non-increasing sequences and $\Gamma_{\min}^{\alpha}$ is the minimum of $\Gamma^{\alpha}$. We must have $R_{k}\geq 0$. Incorporating the results of (D), (D), (D) gives
@@
$+\eta_{\lambda}\left(\alpha^{2}\ell_{21}^{2}\delta_{k}+2\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)r_{k}\right)+\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\left(\xi_{1,k+1}\delta_{k+1}-\xi_{1,k}\delta_{k}\right)+\left(\xi_{2,k+1}r_{k+1}-\xi_{2,k}r_{k}\right)$
$\leq R_{k}-\left(\frac{\eta_{\lambda}}{2}-\ell_{\Gamma}\eta_{\lambda}^{2}-2\xi_{1,k+1}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}-2\xi_{2,k+1}(1+\gamma_{2})\kappa^{2}\eta_{\lambda}^{2}\right)\left\|\nabla\Gamma^{\alpha}(\lambda_{k})\right\|^{2}+\phi_{1}\delta_{k}+\phi_{2}r_{k}$
$+\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\xi_{1,k+1}(1+\gamma_{1}^{-1})\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}+\xi_{2,k+1}(1+\gamma_{2}^{-1})\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}$ (36)
@@
$\phi_{1}$ $=\xi_{1,k+1}\left(4\alpha^{2}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}\ell_{21}^{2}+(1+1/\gamma_{1})\left(1-\frac{\alpha\mu_{2}\eta_{u}}{J}\right)\right)-\xi_{1,k}+2\ell_{\Gamma}\eta_{\lambda}^{2}\alpha^{2}\ell_{21}^{2}+\eta_{\lambda}\alpha^{2}\ell_{21}^{2}$
$+8\xi_{2,k+1}(1+\gamma_{2})\kappa^{2}\eta_{\lambda}^{2}\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)$
$\phi_{2}$ $=\xi_{2,k+1}\left(4(1+\gamma_{2})\kappa^{2}\eta_{\lambda}^{2}\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)+(1+1/\gamma_{2})\left(1-\frac{\alpha\mu_{2}\eta_{w}}{2J}\right)\right)-\xi_{2,k}+4\ell_{\Gamma}\eta_{\lambda}^{2}\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)$
$+2\eta_{\lambda}\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)+8\xi_{1,k+1}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}\alpha^{2}\ell_{21}^{2}.$ (37)
@@
$\phi_{1}$ $=\xi_{1,k+1}\left(\frac{4(1+\gamma_{1})\kappa^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}\ell^{2}+(1+1/\gamma_{1})\left(1-\frac{\mu_{2}\eta_{0}}{JK^{(a-c)}}\right)\right)-\xi_{1,k}+\frac{2\ell_{\Gamma}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}+\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}$
$\quad+\frac{8\xi_{2,k+1}(1+\gamma_{2})\kappa^{2}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}$
$\phi_{2}$ $=\xi_{2,k+1}\left(\frac{4(1+\gamma_{2})\kappa^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}\ell^{2}+(1+1/\gamma_{2})\left(1-\frac{\mu_{2}\eta_{0}}{2JK^{(a-c)}}\right)\right)-\xi_{2,k}+\frac{4\ell_{\Gamma}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2{(b-c)}}}+\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}$
$\quad+\frac{8\xi_{1,k+1}(1+\gamma_{1})\kappa^{2}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2{(b-c)}}}.$ (38)
@@
For $\eta_{0}\leq\frac{8J}{\mu_{2}}$, we have $\frac{\mu_{2}\eta_{0}}{4J}\leq\frac{1}{2}$. Consider that $\xi_{1,k}$ and $\xi_{2,k}$ are non-increasing sequence, then $\xi_{1,k}\geq\xi_{1,k+1}$ and $\xi_{2,k}\geq\xi_{2,k+1}$, we have
$\phi_{1}$ $\leq\xi_{1,k}\left(1+\frac{(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}J}{\mu_{2}\eta_{0}K^{(2b-c-a)}}-\frac{3\mu_{2}\eta_{0}}{4JK^{(a-c)}}\right)-\xi_{1,k}+\frac{2\ell_{\Gamma}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}+\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}+\xi_{2,k}\frac{8J(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}}{\mu_{2}\eta_{0}K^{(2b-c-a)}}\leq 0$
$\phi_{2}$ $\leq\xi_{2,k}\left(1+\frac{(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}J}{\mu_{2}\eta_{0}K^{(2b-c-a)}}-\frac{\mu_{2}\eta_{0}}{4JK^{(a-c)}}\right)-\xi_{2,k}+\frac{4\ell_{\Gamma}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}+\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}+\xi_{1,k}\frac{8J(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}}{\mu_{2}\eta_{0}K^{(2b-c-a)}}\leq 0$

### v2

@@ Appendix A Additional Experiments
A.1 Data Weights across Model Families
Table 7 shows the learned data weights from ScaleBiO for different backbone models under the instruction following setting.
Llama-3-8B Llama-3-13B [^1] https://huggingface.co/Replete-AI/Llama-3-13B Qwen-2-7B Gemma-2-9B GPT-NeoX-20B Yi-34B
source weight source weight source weight source weight source weight source weight
WildChat 0.711 WildChat 0.711 SlimOrca 0.945 Alpaca-pt 0.198 Airoboros 0.986 ShareGPT4 0.627
Airoboros 0.154 ShareGPT4 0.137 LMSYS-Chat 0.008 Alpaca-ko 0.180 ShareGPT4 0.005 Airoboros 0.111
ChatAlpaca 0.119 ChatAlpaca 0.021 ShareGPT4 0.004 Alpaca-it 0.080 ChatAlpaca 0.003 WildChat 0.105
Total 0.984 Total 0.869 Total 0.957 Total 0.458 Total 0.994 Total 0.843
Table 7: Data sources with top-3 weights for LLaMA-3-8B, LLaMA-3-13B, Qwen-2-7B, Gemma-2-9B, GPT-NeoX-20B and Yi-34B in Instruction Following tasks.
Llama-3-8B Qwen-2-7B Gemma-2-9B
source weight source weight source weight
TIGER-Lab/MathInstruct 0.131 TIGER-Lab/MathInstruct 0.132 TIGER-Lab/MathInstruct 0.121
teknium/GPT4-LLM-Cleaned 0.119 ise-uiuc/Magicoder-Evol-Instruct-110K 0.125 DART-Math 0.114
anon8231489123/ShareGPT_Vicuna_unfiltered 0.107 teknium/GPTeacher-General-Instruct 0.102 openbmb/UltraInteract_sft 0.110
Total 0.357 Total 0.359 Total 0.345
Table 8: Data sources with top-3 weights for LLaMA-3-8B, Qwen-2-7B, Gemma-2-9B in Mathematical Reasoning tasks.
A.2 Mathematical Reasoning: Stronger Benchmarks
We conducted additional experiments on mathematical reasoning using stronger benchmarks and a smaller but higher-quality dataset.
Setup
Specifically, we collect 8K prompts uniformly from DART-Math (Tong et al., 2024), Ultra-Interact (Yuan et al., 2024), MathInstruct (Yue et al., 2023), and Orca-Math (Mitra et al., 2024). We then use Deepseek-R1 (DeepSeek-AI et al., 2025) to generate responses with thinking paths to construct question-answer pairs. After obtaining the data, we select 4K training samples using the ScaleBiO method alongside baseline methods and fine-tune the DeepSeek-R1-Distill-Qwen-1.5B model (DeepSeek-AI et al., 2025). The fine-tuned models are then evaluated on the reference sets of AIME24, AIME25, and AIMO25, which contain 30, 30, and 10 questions, respectively.
Method (pass@1) Method (cons@64)
Method AIME 2024 AIME 2025 AIMO 2025 AIME 2024 AIME 2025 AIMO 2025
Uniform 26.7 20.0 10.0 33.3 33.3 30.0
LESS 26.7 20.0 10.0 33.3 36.7 30.0
RHO-LOSS 30.0 20.0 20.0 33.3 33.3 30.0
ScaleBiO 33.3 26.7 20.0 33.3 36.7 30.0
Table 9: Comparison of methods on AIME2024, AIME2025, and AIMO2025 datasets under pass@1 and cons@64 metrics for DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025).
Results
As shown in Table 9, ScaleBiO consistently outperforms the baseline methods across the three benchmarks under the pass@1 metric. For the Cons@64 accuracy, ScaleBiO achieves performance comparable to the baselines. We conjecture that the narrowing gap is due to the limited diversity of the small-sized dataset, which we expect to improve with the inclusion of a larger amount of data. In summary, ScaleBiO demonstrates stable and competitive performance on challenging benchmarks across different evaluation metrics, highlighting the effectiveness of our data selection method.
@@ B.1 Small Scale Experiments
Throughout our small-scale experiments, we use GPT-2 Radford et al. (2019) with 124 million parameters as the backbone model. For bilevel optimization hyperparameters, we set the learning rate to $10^{-2}$ for sampling weights $\lambda$ and $10^{-5}$ for models $u,w$. We run our algorithm for 3 epochs with a batch size of 64 and alpha of 10 while adopting AdamW Loshchilov and Hutter (2017) for optimization.
@@ B.2 Large Scale Experiments
Datasets #Samples Kind License
AlpacaGPT4 (Peng et al., 2023) 52K Instruction Apache-2.0
ShareGPT4 (Chiang et al., 2023) 6K Conversation Apache-2.0
SlimOrca (Lian et al., 2023) 518K Instruction MIT
AlpacaChat (Bian et al., 2023) 20K Conversation Apache-2.0
OpenOrcaGPT4 (Mukherjee et al., 2023) 1M Instruction MIT
WildChat (Zhao et al., 2024) 1M Conversation AI2 ImpACT
LMSYS-Chat (Zheng et al., 2023a) 1M Conversation LMSYS-Chat-1M
GPTeacher ("Teknium", 2023) 89K Instruction MIT
Airoboros (Durbin, 2023) 59K Conversation CC-BY-4.0
@@
Table 10: Training data sources for the Instruction Following task.
Instruction Following
Our training data consists of 18 distinct sources as detailed in Table 10. We collect 9 high-quality datasets and 9 multilingual Alpaca datasets which serve as irrelevant data sources. For each data source, we preprocess by filtering out conversations/instructions that exceed the max length (1024 tokens in our experiments). For our reference dataset $\mathcal{D}_{\text{val}}$ that corresponds to loss $L_{1}$, we prompt GPT4 using the prompt
@@
Here {category} represents one of the 8 categories in MT-Bench and {example} is one example from MT-Bench. In this way, we obtain a reference dataset with 1,200 samples with a similar distribution to MT-Bench. Furthermore, additional 600 samples generated in similar fashions are adopted for hyperparameter tuning for all methods.
Concerning the data reweighting and training process, we first sample 3,000 data from each source for reweighting. Then we sample 10,000 data according to the weights at the end of bilevel optimization to train the backbone model.
For ScaleBiO, the data reweighting process lasts for 3 epochs with $\alpha$ equals to 100 and initial learning rate $10^{-2}$ for weights $\lambda$. The learning rates of models $u,w$ are set to be the same and searched in range $\{10^{-6},2\times 10^{-6},3\times 10^{-6},4\times 10^{-6},5\times 10^{-6},6\times 10^{-6},8\times 10^{-6},10^{-5}\}$. For all the fine-tuning processes, we train the LLM for 1 epoch with an initial learning rate of $8\times 10^{-6}$ and a global batch size of 64. Throughout our experiments, we adopt randomized coordinate descent with AdamW (Pan et al., 2024) and bfloat16 precision for efficient training and inference. Our experiments are conducted on 8 NVIDIA H100 80GB GPUs, where the total computational cost is around $\sim$6K GPU hours. The multi-GPU feature of ScaleBiO is enabled by Pytorch’s FSDP (Zhao et al., 2023).
For baselines, all of them are free to utilize the additional 1,800 MT-Bench-styled samples to ensure a fair comparison with ScaleBiO, where
  • Uniform Weighting: directly sample 10,000 / 18 $\approx$ 5556 samples from each source, along with the additional MT-styled 1,200 samples to conduct supervised fine-tuning.
  • LESS: we stick to settings in its original paper (Xia et al., 2024b), which adopts a warm-up training setup of learning rate $10^{-5}$, batch size 32, maximum sequence length of 1024, number of epochs 4, optimizer Adam with linear decay learning rate schedule. The LoRA setup is also similar, with $r=128$, $\alpha=512$, dropout = 0.1.
  • RHO-LOSS: a training setup of learning rate $10^{-5}$, batch size 32, maximum sequence length 1024, number of epochs 1, optimizer of Adam with cosine decay learning rate schedule. Here the same reference model Qwen-2-1.5B is employed for different settings, which according to the original paper (Mindermann et al., 2022a), is fine given the algorithm’s non-sensitiveness to the reference model.
Mathematical Reasoning
Datasets #Samples Kind License
DART-Math [^11] https://huggingface.co/datasets/hkust-nlp/dart-math-uniform (Tong et al., 2024) 591K Math MIT
SlimOrca [^12] https://huggingface.co/datasets/Open-Orca/SlimOrca (Lian et al., 2023) 518K Instruction MIT
openbmb/UltraInteract_sft [^13] https://huggingface.co/datasets/openbmb/UltraInteract_sft (Yuan et al., 2024) 289K Reasoning MIT
TIGER-Lab/MathInstruct [^14] https://huggingface.co/datasets/TIGER-Lab/MathInstruct (Yue et al., 2023) 262K Reasoning MIT
microsoft/orca-math-word-problems-200k [^15] https://huggingface.co/datasets/microsoft/orca-math-word-problems-200k (Mitra et al., 2024) 200K Math MIT
WizardLMTeam/WizardLM_evol_instruct_V2_196k [^16] https://huggingface.co/datasets/WizardLMTeam/WizardLM_evol_instruct_V2_196k 196K Instruction MIT
ise-uiuc/Magicoder-Evol-Instruct-110K [^17] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K 110K Coding Apache-2.0
anon8231489123/ShareGPT_Vicuna_unfiltered [^18] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered 94K Instruction Apache-2.0
teknium/GPTeacher-General-Instruct [^19] https://huggingface.co/datasets/teknium/GPTeacher-General-Instruct 89K Instruction MIT
teknium/GPT4-LLM-Cleaned [^20] https://huggingface.co/datasets/teknium/GPT4-LLM-Cleaned 55K Instruction Apache-2.0
GSM8K (Cobbe et al., 2021) 7.5K Math MIT
Competition Math [^21] https://huggingface.co/datasets/EleutherAI/hendrycks_math (Hendrycks et al., 2021c) 12.5K Math MIT
bigcode/self-oss-instruct-sc2-exec-filter-50k [^22] https://huggingface.co/datasets/bigcode/self-oss-instruct-sc2-exec-filter-50k 50.7K Coding ODC-By
cais/mmlu [^23] https://huggingface.co/datasets/cais/mmlu (Hendrycks et al., 2020, 2021a) 116K Science MIT
ARC-Easy (Clark et al., 2018) 5.2K Instruction CC-BY-SA-4.0
ARC-Challenge (Clark et al., 2018) 2.6K Instruction CC-BY-SA-4.0
Table 11: Training (above the line) and validation data sources (below the line) for the Mathematical Reasoning task.
The validation set comes from the validation sets (if available) or training sets (if validation is not available) from the validation sources presented in Table 11, where 280 samples are randomly chosen from each source and form a validation set with $280\times 6=1680$ samples. For ScaleBiO, the dataset is proportionally split into two sets with $1,400$ samples and $280$ samples individually, where the former is treated as the $\mathcal{D}_{\text{val}}$ for $L_{1}$ in reweighting and the latter is adopted for hyperparameter tuning. The whole validation set is available to other baselines. Other settings and statistics remain the same as the instruction-following task.
@@ Appendix C Important Lemmas
The results of Lemma 2 can be found in (Kwon et al., 2023) and Lemma B.1 of Chen et al. (2023). From Lemma B.7 in Chen et al. (2023), the following result holds for $\Gamma^{\alpha}(\lambda)$:
@@
The result in Lemma 4 follows from Lemma B.2 of Chen et al. (2023).
@@
    $w_{\ast}^{\alpha}(\lambda)$ is $\ell_{u_{\ast},0}$-Lipschitz continuous where $\ell_{u_{\ast},0}=3\kappa$.
@@
Claim (i) in Lemma 5 can be found in Lemma 2.2 of (Ghadimi and Wang, 2018) and Claim (ii) implies from Lemma 3.2 (setting $\lambda_{1}=\lambda_{2}$) of (Kwon et al., 2023).
@@
Under Assumption 1, if $\alpha>2\ell_{f,1}/\mu_{g}$, then $u_{\ast}(\lambda)$ is $\ell_{\nabla u_{\ast}}$-smooth where $\ell_{\nabla_{\ast}}=\mathcal{O}\left(\frac{\kappa^{2}}{\mu_{2}}\left(\ell_{21}+1\right)\right)$ where the condition number $\kappa=\max\left\{\ell_{10},\ell_{11},\ell_{21},\ell_{22}\right\}/\mu_{2}$
Following Lemma A.3 of (Kwon et al., 2023) and recalling the Lipschitz continuous property of $u_{\ast}(\lambda)$ from Lemma 5, we have this claim is correct.
@@ Appendix D Proofs of Theorem 1
$\mathcal{L}_{D_{k}}^{\alpha}:=L_{1}(\lambda,w;D_{\text{val}}^{k})+\alpha\left(L_{2}(\lambda,w;D_{\text{tr}}^{k})-L_{2}(\lambda,u;D_{\text{tr}}^{k})\right)$ (14)
where $D_{\text{tr}}^{k}$ is i.i.d. from the training dataset $D_{train}$, $D_{\text{val}}^{k}$ are i.i.d. from the validation dataset $D_{val}$ and independent with $D_{\text{train}}$. We use $\mathcal{F}_{k}$ to denote the random information before the iteration $(\lambda_{k},w_{k},u_{k})$, that is $\mathcal{F}_{k}:=\sigma\left(\left\{(\lambda_{k},\omega_{k},u_{k}),D_{k-1},\cdots,D_{1}\right\}\right)$. We use $\mathcal{C}_{k}=\sigma\left(\left\{j_{1},j_{2}\cdots,j_{t-1};r_{1},r_{2},\cdots,r_{t-1}\right\}\right)$ to denote the random information of variables $u,w$ for the randomized block coordinates before the iteration $k$.
@@
$+2(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}\left\|\nabla\Gamma^{\alpha}(\lambda_{k})\right\|^{2}+(1+1/\gamma_{1})\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}$ (31)
@@
$\mathbb{E}\left[\left\|\omega_{\ast}^{\alpha}(\lambda_{k})-\omega_{k+1}\right\|^{2}\mid\mathcal{F}_{k},\mathcal{C}_{k}\right]\leq\left(1-\frac{\alpha\mu_{2}\eta_{w}}{2J}\right)r_{k}+\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}.$ (32)
@@
$R_{k}=\Gamma^{\alpha}(\lambda_{k})-\Gamma_{\min}^{\alpha}+\xi_{k}^{1}\delta_{k}+\xi_{2}^{k}r_{k}$ (34)
where $\xi_{k}^{1},\xi_{k}^{2}>0$ are non-increasing sequences and $\Gamma_{\min}^{\alpha}$ is the minimum of $\Gamma^{\alpha}$. We must have $R_{k}\geq 0$. Incorporating the results of (D), (D), (D) gives
@@
$+\eta_{\lambda}\left(\alpha^{2}\ell_{21}^{2}\delta_{k}+2\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)r_{k}\right)+\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\left(\xi_{k+1}^{1}\delta_{k+1}-\xi_{k}^{1}\delta_{k}\right)+\left(\xi_{k+1}^{2}r_{k+1}-\xi_{k}^{2}r_{k}\right)$
$\leq R_{k}-\left(\frac{\eta_{\lambda}}{2}-\ell_{\Gamma}\eta_{\lambda}^{2}-2\xi_{k+1}^{1}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}-2\xi_{k+1}^{2}(1+\gamma_{2})\kappa^{2}\eta_{\lambda}^{2}\right)\left\|\nabla\Gamma^{\alpha}(\lambda_{k})\right\|^{2}+\phi_{1}\delta_{k}+\phi_{2}r_{k}$
$+\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\xi_{k+1}^{1}(1+\gamma_{1}^{-1})\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}+\xi_{k+1}^{2}(1+\gamma_{2}^{-1})\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}$ (35)
@@
$\phi_{1}$ $=\xi_{k+1}^{1}\left(4\alpha^{2}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}\ell_{21}^{2}+(1+1/\gamma_{1})\left(1-\frac{\alpha\mu_{2}\eta_{u}}{J}\right)\right)-\xi_{k}^{1}+2\ell_{\Gamma}\eta_{\lambda}^{2}\alpha^{2}\ell_{21}^{2}+\eta_{\lambda}\alpha^{2}\ell_{21}^{2}$
$+8\xi_{k+1}^{2}(1+\gamma_{2})\kappa^{2}\eta_{\lambda}^{2}\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)$
$\phi_{2}$ $=\xi_{k+1}^{2}\left(4(1+\gamma_{2})\kappa^{2}\eta_{\lambda}^{2}\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)+(1+1/\gamma_{2})\left(1-\frac{\alpha\mu_{2}\eta_{w}}{2J}\right)\right)-\xi_{k}^{2}+4\ell_{\Gamma}\eta_{\lambda}^{2}\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)$
$+2\eta_{\lambda}\left(\ell_{11}^{2}+\alpha^{2}\ell_{21}^{2}\right)+8\xi_{k+1}^{1}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}\alpha^{2}\ell_{21}^{2}.$ (36)
@@
$\phi_{1}$ $=\xi_{k+1}^{1}\left(\frac{4(1+\gamma_{1})\kappa^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}\ell^{2}+(1+1/\gamma_{1})\left(1-\frac{\mu_{2}\eta_{0}}{JK^{(a-c)}}\right)\right)-\xi_{k}^{1}+\frac{2\ell_{\Gamma}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}+\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}$
$\quad+\frac{8\xi_{k+1}^{2}(1+\gamma_{2})\kappa^{2}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}$
$\phi_{2}$ $=\xi_{k+1}^{2}\left(\frac{4(1+\gamma_{2})\kappa^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}\ell^{2}+(1+1/\gamma_{2})\left(1-\frac{\mu_{2}\eta_{0}}{2JK^{(a-c)}}\right)\right)-\xi_{k}^{2}+\frac{4\ell_{\Gamma}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2{(b-c)}}}+\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}$
$\quad+\frac{8\xi_{k+1}^{1}(1+\gamma_{1})\kappa^{2}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2{(b-c)}}}.$ (37)
@@
For $\eta_{0}\leq\frac{8J}{\mu_{2}}$, we have $\frac{\mu_{2}\eta_{0}}{4J}\leq\frac{1}{2}$. Consider that $\xi_{k}^{1}$ and $\xi_{k}^{2}$ are non-increasing sequence, then $\xi_{k}^{1}\geq\xi_{k+1}^{1}$ and $\xi_{k}^{2}\geq\xi_{k+1}^{2}$, we have
$\phi_{1}$ $\leq\xi_{k}^{1}\left(1+\frac{(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}J}{\mu_{2}\eta_{0}K^{(2b-c-a)}}-\frac{3\mu_{2}\eta_{0}}{4JK^{(a-c)}}\right)-\xi_{k}^{1}+\frac{2\ell_{\Gamma}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}+\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}+\xi_{k}^{2}\frac{8J(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}}{\mu_{2}\eta_{0}K^{(2b-c-a)}}\leq 0$
$\phi_{2}$ $\leq\xi_{k}^{2}\left(1+\frac{(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}J}{\mu_{2}\eta_{0}K^{(2b-c-a)}}-\frac{\mu_{2}\eta_{0}}{4JK^{(a-c)}}\right)-\xi_{k}^{2}+\frac{4\ell_{\Gamma}\ell^{2}(\eta_{\lambda}^{0})^{2}}{K^{2(b-c)}}+\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}+\xi_{k}^{1}\frac{8J(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}}{\mu_{2}\eta_{0}K^{(2b-c-a)}}\leq 0$